# Constraints on Long-Ranged Interactions Between Dark Matter and the Standard Model

### v1

@@ III.3 Dynamical Friction with Net-Charged Dark Matter
Figure 5: The upper bounds on new forces that repel dark matter (DM) particles from one another and either attract or repel dark matter to/from Standard Model (SM) particles, in terms of their range $\lambda$ and their strength relative to gravity $\alpha_{\rm SD}$—see Eq. (1) (with $q_{S}=q_{D}=1$)—in the limit of zero equivalence principle (EP) violation arising from the SM-SM force. Shown are the bounds from combining constraints on DM-DM interactions with constraints on SM-SM interactions we present in Section II, as well as the bounds from dynamical friction in ultrafaint dwarf galaxies (UFDs) that we derive in Section III. The non-dynamical friction bounds are shown for two example DM masses $m_{\rm DM}$, along with a DM mass-independent bound (arising from the need for stars to stay bound to UFDs) that applies only if DM and SM particles repel one another. The dynamical friction bounds are, in this case, independent of DM mass; see the discussion above Eq. (31). We show the dynamical friction bounds for both the attractive and repulsive SM-DM cases, although the distinction between the latter two is small (except that the repulsive case is nonsensical beyond the UFD stellar binding line). To illustrate our sensitivity to the choice of assumed UFD halo profile, the attractive cases are shown for $\gamma=0$ (solid), $\gamma=1$ (dotted), and NFW (dashed) profiles.
Figure 6: The upper bounds on new forces that repel dark matter (DM) particles from one another and either attract or repel dark matter to/from Standard Model (SM) particles, in terms of their range $\lambda$ and their strength relative to gravity $\alpha_{\rm SD}$—see Eq. (1) (with $q_{S}=q_{D}=1$)—in the limit of maximal equivalence principle (EP) violation arising from the SM-SM force. Other than the change in bounds on the SM-SM force in the EP-violating case (which then affect our bounds for reasons discussed around Eqs. (31) and (32)), and the addition of the torsion balance-based test of Milky Way center-directed equivalence principle violation in Ref. [25], this plot is identical to Fig. 5; see that figure’s caption for details.
Several features in Figs. 36 are readily understood. First, note the varying minimal ranges at which any constraint can be set. If not for the modification of inferred halo properties in the presence of new forces, all of the exclusion lines would end at a range of order mpc (depending on the profile shape) because the momentum-transfer cross-section is essentially saturated once every $\chi$ particle within $\lambda$ of a star is scattered by order-one; raising the coupling further has only a logarithmic effect on the heating rate [^2] For simplicity, we conservatively omit this logarithmic effect and omit scattering at ranges greater than $\lambda$ entirely; see Eq. (47) and the discussion that follows it. This minimum range would thus correspond to a total mass of dark matter within $\lambda$ equal to the star’s mass, i.e
@@
Second, it is noticeable that $\gamma=0$ profiles generally give substantially tighter constraints than the $\gamma=1$ and NFW profiles. (In fact, when dark matter attracts itself, $\gamma=1$ and NFW profiles lead to no constraints at all.) This results simply from the much more rapid decrease in dark matter density in the two cuspy profiles than in the cored $\gamma=0$ profile: since the known stellar velocity distribution forces all three profiles to have similar $\rho_{\chi}$ around $R_{*}$, the cuspy profiles have much smaller dark matter densities at larger scale radii [^3] Note that this effect relies significantly on the dark matter density scaling as $\rho_{\chi}\propto r^{-1}$ for radii as small as $R_{*}$. A profile whose density increases as one decreases $r$ below $R_{\chi}$, but whose density levels off at a radius larger than $R_{*}$, would give results somewhere between our cuspy and cored examples.. This effect can be seen for the gravity-only case in Fig. 2. As we discuss in Section III.5, Figs. 910 also show constraints for the $\gamma=1$ and NFW profiles requiring stars to start within $10R_{*}$, instead of within $R_{\chi}$ (for $\gamma=0$ we have $R_{\chi}<10R_{*}$, so this isn’t sensible); this substantially strengthens these constraints, not only because of the reduced deorbiting required, but also because it avoids the very low dark matter densities at large radii.
Most of the other sharp features in these results come from transitions between the various constraints on $\alpha_{\rm DD}$ and $\alpha_{\rm SS}$, due to their downstream effects on the most conservative halo assumptions. The exceptions correspond to situations when one of the halo correction factors above (e.g. when the factors of $1\pm\alpha_{\rm SD}^{\rm(eff)}$ in Eqs. (29)–(32)) change from being essentially equal to $1$ to varying significantly; see, for example, the sudden separation of the SM-DM attracting and repelling lines in Fig. 5 above approximately 30 pc, compared to their essentially identical values at lower ranges.
@@
III.4 Dynamical Friction with Net-Neutral Dark Matter — Plasma Effects
When dark matter is net-neutral—for example, in the case of millicharged dark matter with equal amounts of positive and negative charge—the modifications of inferred halo properties discussed in the previous section no longer apply, since there are no net forces on or by the dark matter collectively. However, dynamical friction can still be significant, since it is independent of the sign of the charge (note the quadratic dependence on $\alpha_{\rm SD}^{2}$ in Eq. (22)).
@@ III.4 Dynamical Friction with Net-Neutral Dark Matter — Plasma Effects
Two complications relative to dynamical friction from gravity alone are present in this case: First, the presence of the dark matter plasma leads to Debye screening of the force, effectively reducing the force’s range. Second, as we noted in Section II, dark matter plasmas may be constrained in the near future by the emergence of two-stream instabilities in various astrophysical systems [11]. As we describe below, dynamical friction cannot set new constraints in this parameter space on its own, but it may be able to in combination with these potential plasma instability constraints.
@@
$\lambda_{D}^{2}=\frac{T_{\chi}}{4\pi(\alpha_{\rm DD}Gm_{\chi}^{2})n_{\chi}}\sim\frac{\sigma_{\chi}^{2}}{4\pi\alpha_{\rm DD}G\rho_{\chi}},$ (34)
@@
This dependence on $\alpha_{\rm DD}$, as opposed to $\alpha_{\rm SD}$, forces us to choose the most conservative—i.e. the largest—allowed $\alpha_{\rm DD}$ when determining whether a particular $\alpha_{\rm SD}$ is excluded, just as we did in some cases in the last section. As we noted previously, however, essentially no constraints on $\alpha_{\rm SD}$ can be set using only existing constraints on $\alpha_{\rm DD}$, as Debye screening renders new forces too short-ranged for dynamical friction to matter. (See also Fig. 10 and the discussion in Section III.5.) We therefore instead present the constraints that would be present if exponentially-growing plasma instabilities in dark matter, as described in Ref. [11], are excluded as well; this leads to non-trivial results, and may be the case in the foreseeable future.
@@
Figure 7: The bounds on new forces when dark matter is net-neutral but stars are not, in terms of their range $\lambda$ and their strength relative to gravity $\alpha_{\rm SD}$—see Eq. (1) (with $q_{S}=q_{D}=1$)—in the limit of zero equivalence principle (EP) violation arising from the SM-SM force. Shown are the existing bounds from combining constraints on DM-DM interactions with constraints on SM-SM interactions we present in Section II, as well as the analogous bounds that would be obtained if exponentially growing plasma instabilities in astrophysical systems were excluded in the future; see Ref. [11]. Also shown (“PI+DF”) are the bounds from dynamical friction in ultrafaint dwarf galaxies (UFDs) that we derive in Section III, again assuming the plasma instability constraints; no meaningful constraints can be obtained without this assumption, due to the effects of Debye screening. The dynamical friction line is for $\gamma=0$ (solid), as no constraint on $\alpha_{\rm SD}$ can be set for the $\gamma=1$ and NFW profiles (but see Section III.5). Two example DM masses $m_{\rm DM}$ are included (though they are not relevant for the plasma instability line alone). For dynamical friction, $m_{\rm DM}$ affects the existing constraints on $\alpha_{\rm DD}$ and thus the maximum strength of Debye screening, although this is only visible for NFW profiles; see Sec. III.4.
Figure 8: The bounds on new forces when dark matter is net-neutral but stars are not, in terms of their range $\lambda$ and their strength relative to gravity $\alpha_{\rm SD}$—see Eq. (1) (with $q_{S}=q_{D}=1$)—in the limit of zero equivalence principle (EP) violation arising from the SM-SM force. Other than the change in bounds on the SM-SM force in the EP-violating case, this plot is identical to Fig. 7; see that figure’s caption for details.
III.5 Constraints on $(\alpha_{\rm SS},\alpha_{\rm DD})$
@@
In Fig. 9, we show the $(\alpha_{\rm SS},\alpha_{\rm DD})$ parameter space for a fixed range of $\lambda=10^{-2}$ pc when all the new forces are attractive. The dark matter mass is chosen to be $m_{\rm DM}=10^{-10}M_{\odot}$, setting the incoherent bound from bullet-cluster in blue, which is the strongest existing constraint on $\alpha_{\rm DD}$ for these parameters; the analogue of this plot for $m_{\rm DM}=1$ GeV would be identical, except that the blue region would be located far above the top of the plot. In the limit of zero equivalence principle violation on the Standard Model side, the strongest constraint on $\alpha_{\rm SS}$ comes from planetary orbits, and is shown in black; constraints on forces with significant equivalence principle violation are off of the left edge of the plot. The green curves illustrate the constraints arising from dynamical friction caused by the new Standard Model-dark matter force for the $\gamma=0$ profile (with our standard assumptions), as well as for an NFW profile with one assumption modified: the initial stellar scale radius is required to be $\leq 10R_{*}$, rather than $\leq R_{\chi}$. (The $\gamma=1$ line with this assumption is essentially identical to this NFW line.) The NFW and $\gamma=1$ curves would be off the plot entirely using our standard, more conservative assumption. The thin, black dotted line, conversely, shows the value of $\alpha_{\rm SD}$ that we can exclude in the $\gamma=0$ case. Note that this significantly understates the exclusion potential of dynamical friction: there is a large, newly-excluded (in the limit of zero equivalence principle violation) region below this line, but this fact is obscured when considering limits on $\alpha_{\rm SD}$ alone.
@@
In Fig. 10, we show the $(\alpha_{\rm SS},\alpha_{\rm DD})$ parameter space for a fixed range of $\lambda=30$ pc when dark matter is net-neutral. Existing constraints from incoherent scattering of dark matter in the bullet cluster assuming $m_{\rm DM}=10^{-10}M_{\odot}$ are shown in blue. Constraints from the known strength of gravity are shown in black, while the red hatched region illustrates where plasma instabilities in astrophysical systems will grow exponentially. The dynamical friction constraints arising from $\alpha_{\rm SD}=\sqrt{\alpha_{\rm SS}\alpha_{\rm DD}}$ are shown in green, including our usual three profile lines and one addition: a dot-dashed line illustrating the (essentially identical) results for $\gamma=1$ and NFW profiles when requiring the initial stellar scale radius to be $\leq 10R_{*}$, rather than $\leq R_{\chi}$. We see that dynamical friction constrains significant new parameter space in this plane for $\gamma=0$ profiles, or with this slightly less conservative assumption. However, when converting this to limits in the $(\lambda,\alpha_{\rm SD})$ plane, we cannot yet set any constraint as there is parameter space at larger $\alpha_{\rm DD}$ and smaller $\alpha_{\rm SS}$, giving the same $\alpha_{\rm SD}$, that is not presently excluded. The hypothetical constraint from plasma instabilities would exclude this large $\alpha_{\rm DD}$ region, allowing dynamical friction with $\gamma=0$ to constrain $\alpha_{\rm SD}$. This is why we only show constraints assuming plasma instabilities are excluded in Section III.4.
@@
IV Conclusion
In this work, we have presented two classes of constraints on long-range forces between Standard Model particles and dark matter. First, constraints on $\alpha_{\rm SD}$ can be obtained by combining constraints on $\alpha_{\rm SS}$ and $\alpha_{\rm DD}$ separately (Eq. (2)), since any new SM-DM force also gives a new SM-SM force and an new DM-DM force. To that end, we have reviewed the strongest existing constraints on these two couplings in order to present the strongest available combined constraints on $\alpha_{\rm SD}$. These combined constraints are, in most models, stronger than any constraints that can be placed directly on $\alpha_{\rm SD}$.

### v2

@@ 3.3 Dynamical Friction with Net-Charged Dark Matter
Figure 5: The upper bounds on new forces that repel dark matter (DM) particles from one another and either attract or repel dark matter to/from Standard Model (SM) particles, in terms of their range $\lambda$ and their strength relative to gravity $\alpha_{\rm SD}$—see Eq. (1.1) (with $q_{S}=q_{D}=1$)—in the limit of zero equivalence principle (EP) violation arising from the SM-SM force. Shown are the bounds from combining constraints on DM-DM interactions with constraints on SM-SM interactions we present in Section 2, as well as the bounds from dynamical friction in ultrafaint dwarf galaxies (UFDs) that we derive in Section 3. The non-dynamical friction bounds are shown for two example DM masses $m_{\rm DM}$, along with a DM mass-independent bound (arising from the need for stars to stay bound to UFDs) that applies only if DM and SM particles repel one another. The dynamical friction bounds are, in this case, independent of DM mass; see the discussion above Eq. (3.25). We show the dynamical friction bounds for both the attractive and repulsive SM-DM cases, although the distinction between the latter two is small (except that the repulsive case is nonsensical beyond the UFD stellar binding line). To illustrate our sensitivity to the choice of assumed UFD halo profile, the attractive cases are shown for $\gamma=0$ (solid), $\gamma=1$ (dotted), and NFW (dashed) profiles.
Figure 6: The upper bounds on new forces that repel dark matter (DM) particles from one another and either attract or repel dark matter to/from Standard Model (SM) particles, in terms of their range $\lambda$ and their strength relative to gravity $\alpha_{\rm SD}$—see Eq. (1.1) (with $q_{S}=q_{D}=1$)—in the limit of maximal equivalence principle (EP) violation arising from the SM-SM force. Other than the change in bounds on the SM-SM force in the EP-violating case (which then affect our bounds for reasons discussed around Eqs. (3.25) and (3.26)), and the addition of the torsion balance-based test of Milky Way center-directed equivalence principle violation in Ref. [27], this plot is identical to Fig. 5; see that figure’s caption for details.
Several features in Figs. 36 are readily understood. First, note the varying minimal ranges at which any constraint can be set. If not for the modification of inferred halo properties in the presence of new forces, all of the exclusion lines would end at a range of order mpc (depending on the profile shape) because the momentum-transfer cross-section is essentially saturated once every $\chi$ particle within $\lambda$ of a star is scattered by order-one; raising the coupling further has only a logarithmic effect on the heating rate [^2] For simplicity, we conservatively omit this logarithmic effect and omit scattering at ranges greater than $\lambda$ entirely; see Eq. (B.6) and the discussion that follows it. This minimum range would thus correspond to a total mass of dark matter within $\lambda$ equal to the star’s mass, i.e
@@
Second, it is noticeable that $\gamma=0$ profiles generally give substantially tighter constraints than the $\gamma=1$ and NFW profiles. (In fact, when dark matter attracts itself, $\gamma=1$ and NFW profiles lead to no constraints at all.) This results simply from the much more rapid decrease in dark matter density in the two cuspy profiles than in the cored $\gamma=0$ profile: since the known stellar velocity distribution forces all three profiles to have similar $\rho_{\chi}$ around $R_{*}$, the cuspy profiles have much smaller dark matter densities at larger scale radii [^3] Note that this effect relies significantly on the dark matter density scaling as $\rho_{\chi}\propto r^{-1}$ for radii as small as $R_{*}$. A profile whose density increases as one decreases $r$ below $R_{\chi}$, but whose density levels off at a radius larger than $R_{*}$, would give results somewhere between our cuspy and cored examples.. This effect can be seen for the gravity-only case in Fig. 2. As we discuss in Section 3.5, Figs. 910 also show constraints for the $\gamma=1$ and NFW profiles requiring stars to start within $10R_{*}$, instead of within $R_{\chi}$ (for $\gamma=0$ we have $R_{\chi}<10R_{*}$, so this isn’t sensible); this substantially strengthens these constraints, not only because of the reduced deorbiting required, but also because it avoids the very low dark matter densities at large radii.
Most of the other sharp features in these results come from transitions between the various constraints on $\alpha_{\rm DD}$ and $\alpha_{\rm SS}$, due to their downstream effects on the most conservative halo assumptions. The exceptions correspond to situations when one of the halo correction factors above (e.g. when the factors of $1\pm\alpha_{\rm SD}^{\rm(eff)}$ in Eqs. (3.23)–(3.26)) change from being essentially equal to $1$ to varying significantly; see, for example, the sudden separation of the SM-DM attracting and repelling lines in Fig. 5 above approximately 30 pc, compared to their essentially identical values at lower ranges.
@@
3.4 Dynamical Friction with Net-Neutral Dark Matter — Plasma Effects
When dark matter is net-neutral—for example, in the case of millicharged dark matter with equal amounts of positive and negative charge—the modifications of inferred halo properties discussed in the previous section no longer apply, since there are no net forces on or by the dark matter collectively. However, dynamical friction can still be significant, since it is independent of the sign of the charge (note the quadratic dependence on $\alpha_{\rm SD}^{2}$ in Eq. (3.16)).
@@ 3.4 Dynamical Friction with Net-Neutral Dark Matter — Plasma Effects
Two complications relative to dynamical friction from gravity alone are present in this case: First, the presence of the dark matter plasma leads to Debye screening of the force, effectively reducing the force’s range. Second, as we noted in Section 2, dark matter plasmas may be constrained in the near future by the emergence of two-stream instabilities in various astrophysical systems [13]. As we describe below, dynamical friction cannot set new constraints in this parameter space on its own, but it may be able to in combination with these potential plasma instability constraints.
@@
$\lambda_{D}^{2}=\frac{T_{\chi}}{4\pi(\alpha_{\rm DD}Gm_{\chi}^{2})n_{\chi}}\sim\frac{\sigma_{\chi}^{2}}{4\pi\alpha_{\rm DD}G\rho_{\chi}},$ (3.28)
@@
This dependence on $\alpha_{\rm DD}$, as opposed to $\alpha_{\rm SD}$, forces us to choose the most conservative—i.e. the largest—allowed $\alpha_{\rm DD}$ when determining whether a particular $\alpha_{\rm SD}$ is excluded, just as we did in some cases in the last section. As we noted previously, however, essentially no constraints on $\alpha_{\rm SD}$ can be set using only existing constraints on $\alpha_{\rm DD}$, as Debye screening renders new forces too short-ranged for dynamical friction to matter. (See also Fig. 10 and the discussion in Section 3.5.) We therefore instead present the constraints that would be present if exponentially-growing plasma instabilities in dark matter, as described in Ref. [13], are excluded as well; this leads to non-trivial results, and may be the case in the foreseeable future.
@@
Figure 7: The bounds on new forces when dark matter is net-neutral but stars are not, in terms of their range $\lambda$ and their strength relative to gravity $\alpha_{\rm SD}$—see Eq. (1.1) (with $q_{S}=q_{D}=1$)—in the limit of zero equivalence principle (EP) violation arising from the SM-SM force. Shown are the existing bounds from combining constraints on DM-DM interactions with constraints on SM-SM interactions we present in Section 2, as well as the analogous bounds that would be obtained if exponentially growing plasma instabilities in astrophysical systems were excluded in the future; see Ref. [13]. Also shown (“PI+DF”) are the bounds from dynamical friction in ultrafaint dwarf galaxies (UFDs) that we derive in Section 3, again assuming the plasma instability constraints; no meaningful constraints can be obtained without this assumption, due to the effects of Debye screening. The dynamical friction line is for $\gamma=0$ (solid), as no constraint on $\alpha_{\rm SD}$ can be set for the $\gamma=1$ and NFW profiles (but see Section 3.5). Two example DM masses $m_{\rm DM}$ are included (though they are not relevant for the plasma instability line alone). For dynamical friction, $m_{\rm DM}$ affects the existing constraints on $\alpha_{\rm DD}$ and thus the maximum strength of Debye screening, although this is only visible for NFW profiles; see Sec. 3.4.
Figure 8: The bounds on new forces when dark matter is net-neutral but stars are not, in terms of their range $\lambda$ and their strength relative to gravity $\alpha_{\rm SD}$—see Eq. (1.1) (with $q_{S}=q_{D}=1$)—in the limit of zero equivalence principle (EP) violation arising from the SM-SM force. Other than the change in bounds on the SM-SM force in the EP-violating case, this plot is identical to Fig. 7; see that figure’s caption for details.
3.5 Constraints on $(\alpha_{\rm SS},\alpha_{\rm DD})$
@@
In Fig. 9, we show the $(\alpha_{\rm SS},\alpha_{\rm DD})$ parameter space for a fixed range of $\lambda=10^{-2}$ pc when all the new forces are attractive. The dark matter mass is chosen to be $m_{\rm DM}=10^{-10}M_{\odot}$, setting the incoherent bound from bullet-cluster in blue, which is the strongest existing constraint on $\alpha_{\rm DD}$ for these parameters; the analogue of this plot for $m_{\rm DM}=1$ GeV would be identical, except that the blue region would be located far above the top of the plot. In the limit of zero equivalence principle violation on the Standard Model side, the strongest constraint on $\alpha_{\rm SS}$ comes from planetary orbits, and is shown in black; constraints on forces with significant equivalence principle violation are off of the left edge of the plot. The green curves illustrate the constraints arising from dynamical friction caused by the new Standard Model-dark matter force for the $\gamma=0$ profile (with our standard assumptions), as well as for an NFW profile with one assumption modified: the initial stellar scale radius is required to be $\leq 10~{}R_{*}$, rather than $\leq R_{\chi}$. (The $\gamma=1$ line with this assumption is essentially identical to this NFW line.) The NFW and $\gamma=1$ curves would be off the plot entirely using our standard, more conservative assumption. The thin, black dotted line, conversely, shows the value of $\alpha_{\rm SD}$ that we can exclude in the $\gamma=0$ case. Note that this significantly understates the exclusion potential of dynamical friction: there is a large, newly-excluded (in the limit of zero equivalence principle violation) region below this line, but this fact is obscured when considering limits on $\alpha_{\rm SD}$ alone.
@@
In Fig. 10, we show the $(\alpha_{\rm SS},\alpha_{\rm DD})$ parameter space for a fixed range of $\lambda=30$ pc when dark matter is net-neutral. Existing constraints from incoherent scattering of dark matter in the bullet cluster assuming $m_{\rm DM}=10^{-10}M_{\odot}$ are shown in blue. Constraints from the known strength of gravity are shown in black, while the red hatched region illustrates where plasma instabilities in astrophysical systems will grow exponentially. The dynamical friction constraints arising from $\alpha_{\rm SD}=\sqrt{\alpha_{\rm SS}\alpha_{\rm DD}}$ are shown in green, including our usual three profile lines and one addition: a dot-dashed line illustrating the (essentially identical) results for $\gamma=1$ and NFW profiles when requiring the initial stellar scale radius to be $\leq 10R_{*}$, rather than $\leq R_{\chi}$. We see that dynamical friction constrains significant new parameter space in this plane for $\gamma=0$ profiles, or with this slightly less conservative assumption. However, when converting this to limits in the $(\lambda,\alpha_{\rm SD})$ plane, we cannot yet set any constraint as there is parameter space at larger $\alpha_{\rm DD}$ and smaller $\alpha_{\rm SS}$, giving the same $\alpha_{\rm SD}$, that is not presently excluded. The hypothetical constraint from plasma instabilities would exclude this large $\alpha_{\rm DD}$ region, allowing dynamical friction with $\gamma=0$ to constrain $\alpha_{\rm SD}$. This is why we only show constraints assuming plasma instabilities are excluded in Section 3.4.
@@
4 Conclusion
In this work, we have presented two classes of constraints on long-range forces between Standard Model particles and dark matter. First, constraints on $\alpha_{\rm SD}$ can be obtained by combining constraints on $\alpha_{\rm SS}$ and $\alpha_{\rm DD}$ separately (Eq. (2.1)), since any new SM-DM force also gives a new SM-SM force and an new DM-DM force. To that end, we have reviewed the strongest existing constraints on these two couplings in order to present the strongest available combined constraints on $\alpha_{\rm SD}$. These combined constraints are, in most models, stronger than any constraints that can be placed directly on $\alpha_{\rm SD}$.